ENERGY TRANSFER PROCESSES BETWEEN DIATOMIC MOLECULES IN OPTICALLY PUMPED PLASMAS ^{a}

\underline{E. PLÖNJES ^{b}}, P. PALM^{b}, I. ADAMOVICH, J.W. RICH, Department of Mechanical Engineering, The Ohio State University, Columbus, OH 43210.

Energy transfer processes in collision between highly excited diatomic molecules and ionization of the molecules in energy pooling collisions are studied. In a CO-Ar, CO-Ar-N_{2}, or CO-Ar-O_{2} gas mixture carbon monoxide is excited into low vibrational states v = 1, ..., 10 by absorption of CO laser radiation in an optical absorption cell. Vibration-vibration exchange pumping results in the diffusion of energy into very high vibrational states up to level v = 40. Also vibrational energy is transferred to the vibrational modes of N_{2} and O_{2}. Ionization in optically pumped CO occurs by associative ionization mechanism in collisions of two excited CO molecules when their total vibrational energy exceeds the ionization energy. The effect of the electrons created in the plasma on the vibrational distribution function of CO and the influence of N_{2} and O_{2} is studied by FTIR spectroscopy on the Δv = 2 emission. Transfer of vibrational energy to electronic states is monitored using vacuum UV spectroscopy on the 4th positive emission of CO. Electron densities and ionization-rate coefficients are measured in a non-self-sustained, low electric field Thomson discharge. A kinetic model incorporating a coupled master equation for the CO, N_{2}, and O_{2} vibrational level populations, and the Boltzmann equation for the electrons is compared to the experimental results.

aResearch is primarily funded by AFOSR and DDRE under the Air Plasma Rampart MURI and AFOSR Space Propulsion and Power Program under Grant No. F49620-96-1-0184

bVisiting Scholar at The Ohio State University, permanently at the Institut für Angewandte Physik der Universität Bonn, 53115 Bonn, Germany